# Few-Shot Learning with Uncertainty-based Quadruplet Selection for Interference Classification in GNSS Data

### v1

@@ Abstract
publicationid: pubid: 978-1-6654-0575-1/22/\$31.00 ©2022 IEEE
@@ I Introduction
The remainder of this paper is organized as follows. Section II summarizes related work for GNSS-based classification and Section III gives a background on FSL and pairwise learning. In Section IV, we propose our uncertainty-based quadruplet selection method for FSL. We show details of our recorded dataset in Section V. Section VI summarizes evaluation results and we conclude in Section VII.
@@ IV-B Pairwise Learning
(a) Class 0.
(b) Class 1.
(c) Class 2.
(d) Class 3.
(e) Class 4.
(f) Class 5.
(g) Class 6.
(h) Class 7.
(i) Class 8.
(j) Class 9.
@@
Figure 2: Exemplary spectrogram samples without interference (classes 0 to 2) and with interference (classes 3 to 10) between intensity [0, -90] with logarithmic scale. The x-axis shows the time in ms. The y-axis shows the freqeuncy in MHz.
(a) Contrastive loss.
(b) Triplet loss.
@@ VI Evaluation
For all experiments, we use Nvidia Tesla V100-SXM2 GPUs with 32 GB VRAM equipped with Core Xeon CPUs and 192 GB RAM. We use the vanilla Adam optimizer with a learning rate of $0.01$, decay of $0.0005$, and train the standard models for 100 epochs and the models with pairwise learning for 200 epochs. First, we present baseline results of FSL and for different layer adaptations. Next, we propose an evaluation of the uncertainty computation and our triplet and quadruplet loss with hyperparameter searches. In general, we train all possible class permutations for pre-training, adaptation, and evaluation, but present only results for the specific permutation (train classes: 0, 1, 2, 4, 5, 8, and 11; adaptation classes: 3, 7, 9, and 10). However, our statements can be generalized to all permutations. We train each method $M=10$ times and present the mean and standard variance F2-score.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
(a) Confusion matrix.
(b) Aleatoric uncertainty.
@@
Figure 6: Confusion matrices and uncertainties for the PN (left) and the PN with quadruplet loss (right).
@@ Acknowledgments
This work has been carried out within the DARCY project, funding code 50NA2017, sponsored by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), the German Aerospace Center (DLR), the Bundesnetzagentur (BNetzA), and the Federal Agency for Cartography and Geodesy (BKG).

### v2

@@ Abstract
https://gitlab.cc-asp.fraunhofer.de/darcy_gnss/FIOT_highway
@@
publicationid: pubid: 979-8-3503-8078-1/24/\$31.00 ©2024 IEEE
@@ I Introduction
Section II summarizes related work and Section III gives a background on FSL and pairwise learning. In Section IV, we propose our quadruplet selection method for FSL. We show details of our dataset in Section V. Section VI summarizes evaluation results and we conclude in Section VII.
@@ IV-B Pairwise Learning
Figure 2: Exemplary spectrogram samples without interference (classes 0 to 2) and with interference (classes 3 to 10) between intensity [0, -90] with logarithmic scale. The x-axis shows the time in ms. The y-axis shows the frequency in MHz.
@@ VI Evaluation
For all experiments, we use Nvidia Tesla V100-SXM2 GPUs with 32 GB VRAM equipped with Core Xeon CPUs and 192 GB RAM. We use the vanilla SGD optimizer with a learning rate of $0.01$, decay of $0.0005$, and train the standard models for 100 epochs and the models with pairwise learning for 200 epochs. First, we present baseline results of FSL and for different layer adaptations. Next, we propose an evaluation of the uncertainty computation and our triplet and quadruplet loss with hyperparameter searches. In general, we train all possible class permutations for pre-training, adaptation, and evaluation, but present only results for the specific permutation (train classes: 0, 1, 2, 4, 5, 8, and 11; adaptation classes: 3, 7, 9, and 10). However, our statements can be generalized to all permutations. We train each method $M=10$ times and present the mean and standard variance F2-score.
@@
Figure 6: Confusion matrices and uncertainties for the PN (left) and the PN with quadruplet loss (right) averaged over 10 runs.
@@ Acknowledgments
This work has been carried out within the DARCII project, funding code 50NA2401, sponsored by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) and supported by the German Aerospace Center (DLR), the Bundesnetzagentur (BNetzA), and the Federal Agency for Cartography and Geodesy (BKG).